A student measures the molar solubility of magnesium fluoride in a water solution to be 1.19×10-3 M.
Based on her data, the solubility product constant for this compound is__________
Answer #1

At equilibrium:

MgF2 <----> Mg2+ + 2 F-

   s 2s

Ksp = [Mg2+][F-]^2

Ksp = (s)*(2s)^2

Ksp = 4(s)^3

Ksp = 4(1.19*10^-3)^3

Ksp = 6.741*10^-9

Answer: 6.74*10^-9